OpenStudy (anonymous):

~~~~~~~~~~FAN~~~~~~~~~~~~~MEDAL~~~~~~~~ !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (anonymous):

Did you just type something but it didnt show?

HanAkoSolo (jamierox4ev3r):

2x^2+13x+15 (2*15=30) (x+10)(x+3)

HanAkoSolo (jamierox4ev3r):

no thats the first factorization for the first choice, i just wasn't done yet ^_^

HanAkoSolo (jamierox4ev3r):

based from this can u do the other one?

HanAkoSolo (jamierox4ev3r):

@SOS101

OpenStudy (anonymous):

Sorry my bro took my laptop

OpenStudy (anonymous):

I really have a hard time with these

HanAkoSolo (jamierox4ev3r):

lol ok :P i will help u if u still need it

HanAkoSolo (jamierox4ev3r):

the second one: 2x^2+27x+13 (2*13) (x+26)(x+1) THese are the two factorizations, all you have to do is multiply the first number by the last and make sure that the factor goes in to the middle number (26*1=26, and 26+1=27)

HanAkoSolo (jamierox4ev3r):

if you need more explanation, just mention my name. Also, srry it took me a while to respond, I was briefly AFK

HanAkoSolo (jamierox4ev3r):

@SOS101

OpenStudy (anonymous):

Thank You so much , sorry for my tardiness @Jamierox4ev3r

OpenStudy (anonymous):

@Jamierox4ev3r I don't think your factorizations are correct: (x+10)(x+3)=x^2 + 13x + 30, not 2x^2 + 13x + 15 (x+26)(x+1)=x^2 + 27x + 26, not 2x^2 + 27x + 13

OpenStudy (anonymous):

@vinnv226 could you explain in a more detailed way how you got those answers

OpenStudy (anonymous):

I just foiled out the answers that Jamie gave, and showed that they were not equal to the original polynomial

OpenStudy (anonymous):

I'm not good at foiling can you show the way that you foiled it @vinnv226